OpenStudy (anonymous):

Solve by Substitution: 2x + y = 0 x – 3y = -7

OpenStudy (anonymous):

2x+y=0 -2(x-3y=7) have : 2x+y=0 -2x+6y=14 Now sum : -2x+2x+6y+y=14+0 have : 7y=12 y=2 Understood ?:)

OpenStudy (anonymous):

And for find x : 2x+y=0 2x+2=0 x=1

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Welcome boy .